Find the total energy of a moving object of mass 6.3 π‘˜π‘” and momentum 11.5 Γ— 106 π‘˜π‘”π‘šπ‘  βˆ’1 .Β 


1
Expert's answer
2021-12-07T20:12:50-0500
"E=\frac{p^2}{2m}\\E=\frac{(11.5\cdot10^6)^2}{2(6.3)}=1.05\cdot10^{13}\ J"